find the equation of the line parallel to 2x+5y=10, through the point (15,-4).Write the answer in slope-intercept form. Write your answer using integers or simplified fractions
Answer by stanbon(75887)

find the equation of the line parallel to 2x+5y=10, through the point (15,-4).Write the answer in slope-intercept form. Write your answer using integers or simplified fractions
------
Find the slope of 2x+5y = 10
5y = -2x+10
y = (-2/5)x + 2
----
slope = -2/5
==================
Equation Form: y = mx + b
You are y = -4 when x = 15, and you know m = -2/5
----
Solve for "b":
-4 = (-2/5)15 + b
-4 = -6 + b
b = 2
----
Equation:
y = (-2/5)x + 2
====================
Cheers,
Stan H.
=======================